# Continual Learning in Medical Image Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v3

@@ 2.4 Continual learning pipeline
The design of a CL pipeline is illustrated in  Figure 1 which provides an overview of the key components and stages involved in the construction of the CL pipeline. Given a problem statement, first, we need to identify which CL scenarios it falls under, i.e., whether there is a possibility of domain shifts in future data, inclusion of new classes, or the end application, i.e., the task itself may change. For example, if we want to develop a breast cancer classification model to be able to work on the H&E dataset from different centers then the datasets across centers may have drift and hence fall into the domain incremental scenario of CL. Detailed information about CL scenarios is provided in Section 3. Once we have identified the CL scenario, training, and testing, datasets need to be prepared to mimic a continual stream of datasets arriving one after another. The sequence of datasets is frequently referred to as tasks or episodes in literature. We also use these two terms interchangeably in this manuscript. For each episode, separate training and testing data is prepared; thus, the pipeline requires four train-test pairs to develop and evaluate a CL model for four sequential datasets. Once the datasets are ready, a CL strategy suitable to the task at hand is identified and deployed in any off-the-shelf deep classification or segmentation model. There are various CL strategies, some offering privacy-preserved learning, while some offer better performance but at the cost of more resources, storage, and privacy violations as they store some past data. Typically, the model is trained on the first training episode and evaluated on testing data from all the episodes. After this, the training shifts to the next episode, where the inclusion of partial training data from the previous episode is possible. Here, updating the model with new training data, the evaluation is again done on all the testing data, and the process repeats until the last episode. Application-specific performance metrics (e.g., accuracy, dice similarity coefficient, etc.) computed on testing data of each episode are observed and analyzed over the sequence. Then, we can compute various metrics on top of it to quantify forgetting and forward transfer. Lastly, the CL framework is evaluated against state-of-the-art works and non-CL methods offering upper and lower bounds of performance. Joint or cumulative training gives the highest average performance, whereas naively finetuning on the current episode gives the lowest performance Kaustaban et al. (2022); Lenga et al. (2020).
@@
Figure 3: CL-based research contributions over the year
@@ 2.5 Application of continual learning
Disease progression modeling: CL is crucial in tracking the progression of diseases such as Alzheimer’s Disease (AD). Models must adapt to new patient data over time, incorporating the latest diagnostic criteria and treatment strategies Li et al. (2020). Recent work was proposed for modeling AD progression in a CL manner that respects the longitudinal data sets coming in sequence and ensures equal prediction accuracy for future visits Zhang and Wang (2019).
@@ 3 Continual learning scenarios
Depending upon what kind of change in the sequence of data is expected, the existing CL scenarios can be broadly categorized into five categories, viz., (a) data-incremental, (b) class-incremental, (c) task-incremental, (d) domain-incremental, and (e) hybrid. Figure 4 shows the percentages of the above-mentioned incremental CL scenarios exploited in recent years. The statistics reflect that domain-incremental is highly explored with a 51.9% ratio out of five major settings, followed by the class-incremental settings having 26.9% works and data-incremental being the least challenging and rarely explored for different medical image analysis applications.
@@
Kaustaban et al. (2022) Histopathology data-based tumor classification (4 Ep.) Ep. was created in instance incremental fashion from colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes, (i.e., all 9 classes are present in each episode)
@@ 3.1 Instance incremental scenarios
Instance Incremental Scenario (IIS), also termed as data incremental scenario, involves the continuous learning process where the model keeps adapting to incoming data streams that come from the same data distribution. This scenario doesn’t involve dealing with entirely new categories or significant changes in data patterns, making it generally the least challenging among all other CL scenarios. In the medical domain, especially digital pathology, a common practice is for expert pathologists to annotate a dataset in multiple stages or batches Kaustaban et al. (2022). Each of these annotation batches can be thought of as representing samples from the same underlying data distribution, ensuring some consistency in the data source across different batches. Thus, IIS scenarios deal with the condition where the samples of a dataset are sequentially made available. Owning to its simplicity, this category is not very popular. Only a few classification applications are explored in this scenario, as mentioned in Table 2. For example, in the breast cancer classification application, Chen and Tang (2022) set up IIS on the BreakHis Spanhol et al. (2015) dataset. More and more samples from each of the malignant and benign classes are added over the subsequent episodes. Ravishankar et al. (2019) try to exhibit a scenario for pneumothorax detection in a hospital where data arrives in an incremental fashion. They sequentially fed 2000 samples in 4 subsequent episodes from an X-ray dataset for Pneumothorax identification consisting of a total of 8000 samples.
@@
Li et al. (2020) Dermoscopic images based skin disease classification (4-20 Ep.) (a) CIS for Skin8 dataset having total 8 classes on which 2 classes per episode used, (b) CIS for Skin40 dataset from ISIC2019 (Tschandl et al. 2018) consists of 40 classes of skin disease images collected from the internet, here 2/5/10 classes per episode used in 3 exp, (c) non-medical dataset CIFAR100 with 100 classes, here 5/10/20 classes per episode used in 3 exp
Liu et al. (2022) Class incremental segmentation of abdomen organs (4 Ep.) 4 Ep. created for incremental segmentation of liver, spleen, pancreas, right kidney, and left kidney organs from CT datasets (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
Derakhshani et al. (2022) CIS: disease classification (4 Ep.) 4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (taken from MedMNIST repository (Yang et al. 2023a))
@@
Kaustaban et al. (2022) Histopathology data based tumor classification (4 Ep.) 4 Ep. were created in CIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total 9 classes
Chen and Tang (2022) Histopathology data-based breast cancer classification (4 Ep.) 4 Ep. created in CIS from BreakHis dataset
@@
Chee et al. (2023) Cancer classification (3,5 Ep.), diabetic retinopathy classification (2 Ep.), skin lesions classification (3,4 Ep.) (a) Histopathology data based human colorectal cancer classification: CCH5000 (Kather et al. 2016) (total 8 classes, Exp=(4,1,1,1,1), Exp=(4,2,2)), (b) diabetic retinopathy (DR) classification using retinal images: EyePACS (kag ) (total 5 classes: no DR, mild DR, moderate DR, severe DR and proliferative DR, Exp=(3,2)), (c) pigmented skin lesions classification: HAM10000 (Tschandl et al. 2018) (total 7 classes: Exp=(4,1,1,1), Exp=(4,2,1))
Zhang et al. (2023c) Disease classification in CIS (4-10 Ep.) (a) skin lesions classification with Skin8 dataset (International Skin Imaging Collaboration (ISIC) (Tschandl et al. 2018)) having 8 classes distributed over 4 Ep., (b) Path16, a pathology image collection from various public sources: total 16 classes distributed over 7 episodes, (c) CIFAR100: non-medical images, 100 classes distributed over 5-10 episodes
@@
Sadafi et al. (2023) Microscopic images based WBC classification (4 Ep.) With 3 different datasets, 3 separate class-incremental scenarios each having 4 episodes: (a) (3,3,3,4) classes from Matek-19 (Matek et al. 2019), (b) (3,3,3,4) classes from INT-20, and (c) (3,2,2,3) classes from Acevedo-20 (Acevedo et al. 2020)
@@ 3.2 Class incremental scenarios
Class Incremental Scenario (CIS) refers to a situation where the aim is to adapt the model to accommodate novel classes from new data streams. Given that each batch of these unseen classes offers knowledge that largely differs from the previous batches, this scenario is anticipated to be the most challenging among continual learning scenarios. An instance of this challenge is incorporating new types of tissues into a tumor classification model or introducing novel cell types into a model designed for AI-assisted diagnosis in cell detection. In this scenario, there is only one incremental task that gets additions of new classes in subsequent episodes. Some works assume mutually exclusive classes, whereas some overlapping classes.
@@
Kaustaban et al. (2022) Histopathology data based tumor classification (4 Ep.) 4 Ep. were created in TIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes
Derakhshani et al. (2022) Disease classification (4 Ep.) 4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (MedMNIST repository (Yang et al. 2023a))
Bera et al. (2023) Three tasks involving different organs for three different binary segmentation tasks (3 Ep.) MRI data for binary prostate, spleen, hippocampus segmentation: Promise12 (Litjens et al. 2014) (prostate) $\rightarrow$ MSD (NCI ) (spleen) $\rightarrow$ Drayd (Denovellis et al. 2021) (hippocampus)

### v4

@@ 2.4 Continual learning pipeline
The design of a CL pipeline is illustrated in  Figure 1 which provides an overview of the key components and stages involved in the construction of the CL pipeline. Given a problem statement, first, we need to identify which CL scenarios it falls under, i.e., whether there is a possibility of domain shifts in future data, inclusion of new classes, or the end application, i.e., the task itself may change. For example, if we want to develop a breast cancer classification model to be able to work on the H&E dataset from different centers then the datasets across centers may have drift and hence fall into the domain incremental scenario of CL. Detailed information about CL scenarios is provided in Section 3. Once we have identified the CL scenario, training, and testing, datasets need to be prepared to mimic a continual stream of datasets arriving one after another. The sequence of datasets is frequently referred to as tasks, experiences, or episodes in literature. We also use these terms interchangeably in this manuscript. For each episode, separate training and testing data is prepared; thus, for a given sequence of four datasets, the pipeline requires four train-test pairs to develop and evaluate a CL model. Once the datasets are ready, a CL strategy suitable to the application at hand is identified and deployed in any off-the-shelf deep classification or segmentation model. There are various CL strategies, some offering privacy-preserved learning, while some offer better performance but at the cost of more resources, storage, and privacy violations as they store some past data. Typically, the model is trained on the first training episode and evaluated on testing data from all the episodes. After this, the training shifts to the next episode, where the inclusion of partial training data from the previous episode is possible. Here, updating the model with new training data, the evaluation is again done on all the testing data, and the process repeats until the last episode. Application-specific performance metrics (e.g., accuracy, dice similarity coefficient, etc.) computed on testing data of each episode are observed and analyzed over the sequence. Then, we can compute various metrics on top of it to quantify forgetting and forward transfer. Lastly, the CL framework is evaluated against state-of-the-art works and non-CL methods offering upper and lower bounds of performance. Joint or cumulative training gives the highest average performance, whereas naively finetuning on the current episode gives the lowest performance Kaustaban et al. (2022); Lenga et al. (2020).
@@
Figure 3: CL-based research contributions over the years. Percentages represent the number of CL papers in the medical domain each year, showing the increasing trend and growing importance of CL research.
@@ 2.5 Application of continual learning
Disease progression modeling: CL is crucial in tracking the progression of diseases such as Alzheimer’s Disease (AD). Models must adapt to new patient data over time, incorporating the latest diagnostic criteria and treatment strategies Li et al. (2020a). Recent work was proposed for modeling AD progression in a CL manner that respects the longitudinal data sets coming in sequence and ensures equal prediction accuracy for future visits Zhang and Wang (2019a).
@@ 3 Continual learning scenarios
Depending upon what kind of change in the sequence of data is expected, the existing CL scenarios can be broadly categorized into five categories, viz., (a) instance-incremental, (b) class-incremental, (c) task-incremental, (d) domain-incremental, and (e) hybrid. Figure 4 shows the percentages of the above-mentioned incremental CL scenarios exploited in recent years. The statistics reflect that domain-incremental is highly explored with a 44.4% ratio out of five major settings, followed by the class-incremental settings having 32.1% works and instance-incremental being the least challenging and rarely explored for different medical image analysis applications.
@@
Kaustaban et al. (2022) Histopathology data-based tumor classification (4 Ep.) Tasks were created in instance incremental fashion from colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes, (i.e., all 9 classes are present in each episode)
Wei et al. (2023) Brain tumor segmentation (8 Ep.) Starting from 160 samples, authors incrementally add 40 more samples after each training session from the LGG Segmentation dataset (Buda et al. 2019)
Bringas et al. (2024) Alzheimer’s disease stage identification using motion-sensor (2 Ep.,3 Ep.,4 Ep.) 3 experiments were designed with 2, 3, and 4 episodes having per-episode 374, 249, and 187 samples, respectively.
@@ 3.1 Instance incremental scenarios
Instance Incremental Scenario (IIS), also termed as data incremental scenario, involves the continuous learning process where the model keeps adapting to incoming data streams that come from the same data distribution. This scenario doesn’t involve dealing with entirely new categories or significant changes in data patterns, making it generally the least challenging among all other CL scenarios. In the medical domain, especially digital pathology, a common practice is for expert pathologists to annotate a dataset in multiple stages or batches Kaustaban et al. (2022). Each of these annotation batches can be thought of as representing samples from the same underlying data distribution, ensuring some consistency in the data source across different batches. Thus, IIS scenarios deal with the condition where the samples of a dataset are sequentially made available. Owning to its simplicity, this category is not very popular. Only a few classification applications are explored in this scenario, as mentioned in Table 3. For example, in the breast cancer classification application, Chen and Tang (2022) set up IIS on the BreakHis Spanhol et al. (2015) dataset. More and more samples from each of the malignant and benign classes are added over the subsequent episodes. Ravishankar et al. (2019) try to exhibit a scenario for pneumothorax detection in a hospital where data arrives in an incremental fashion. They sequentially fed 2000 samples in 4 subsequent episodes from an X-ray dataset for Pneumothorax identification consisting of a total of 8000 samples. A label drift is also possible in real-world IIS settings. However, the majority of previous works deliberately curate a similar amount of samples for a particular class over episodes, and hence, label drift phenomena are not considered. One potential solution to produce the desired number of samples for classes affected by label-shift from past episodes would be to use a generator model such as GAN Byun et al. (2023).
@@
Li et al. (2020a) Dermoscopic images based skin disease classification (4-20 Ep.) (a) CIS for Skin8 dataset having total 8 classes on which 2 classes per episode used, (b) CIS for Skin40 dataset from ISIC2019 (Tschandl et al. 2018) consists of 40 classes of skin disease images collected from the internet, here 2/5/10 classes per episode used in 3 experiments, (c) non-medical dataset CIFAR100 with 100 classes, here 5/10/20 classes per episode used in 3 experiments.
Liu et al. (2022) Class incremental segmentation of abdomen organs (4 Ep.) 4 episodes were created for incremental segmentation of liver, spleen, pancreas, right kidney, and left kidney organs from CT datasets (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
Derakhshani et al. (2022) Disease classification (4 Ep.) TissueMNIST, OrganaMNIST, PathMNIST, and BloodMNIST datasets from MedMNIST repository (Yang et al. 2023a) were used in 4 experiments, each with 4 episodes.
@@
Kaustaban et al. (2022) Histopathology data based tumor classification (4 Ep.) 4 episodes were created in CIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total 9 classes
Chen and Tang (2022) Histopathology data-based breast cancer classification (4 Ep.) 4 episodes created in CIS from BreakHis dataset
@@
Chee et al. (2023) (a) Cancer classification (3,5 Ep.), (b) diabetic retinopathy classification (2 Ep.), (c) skin lesions classification (3,4 Ep.) (a) Histopathology data based human colorectal cancer classification: CCH5000 (Kather et al. 2016) (total 8 classes, Exp=(4,1,1,1,1) classes, Exp=(4,2,2) classes), (b) diabetic retinopathy (DR) classification using retinal images: EyePACS (kag ) (total 5 classes: no DR, mild DR, moderate DR, severe DR and proliferative DR, Exp=(3,2) classes), (c) pigmented skin lesions classification: HAM10000 (Tschandl et al. 2018) (total 7 classes: Exp=(4,1,1,1) classes, Exp=(4,2,1) classes)
Zhang et al. (2023c) Disease classification (4-10 Ep.) (a) skin lesions classification with Skin8 dataset (International Skin Imaging Collaboration (ISIC) (Tschandl et al. 2018)) having 8 classes distributed over 4 episodes, (b) Path16, a pathology image collection from various public sources: total 16 classes distributed over 7 episodes, (c) CIFAR100: non-medical images, 100 classes distributed over 5-10 episodes
@@
Sadafi et al. (2023) Microscopic images based WBC classification (4 Ep.) With 3 different datasets, 3 separate CIS experiments each having 4 episodes: (a) (3,3,3,4) classes from Matek-19 (Matek et al. 2019), (b) (3,3,3,4) classes from INT-20, and (c) (3,2,2,3) classes from Acevedo-20 (Acevedo et al. 2020)
Wang et al. (2023b) (a) segmentation for endoscopy (2-3 Ep.), (b) surgical instrument segmentation (2 Ep.) Incrementally adding segmentation structure from EDD2020 (Ali et al. 2021, 2020) dataset which includes 5 classes as Barrett’s esophagus, cancer, high-grade dysplasia, polyp, and suspicious. Initially, 3-4 classes are in 1st episode, and then 2 or 1 classes are incrementally added in subsequent episodes. (b) EndoVis18 (Allan et al. 2020) and EndoVis17 (Allan et al. 2019) datasets as two episodes having some overlapping classes
Hua et al. (2023) sEMG-based gesture classification (4 Ep.) 2 experiments were created, each with 4 episodes using Ninapro DB2 sEMG dataset for gesture recognition. The first experiment has (8, 11, 14, 17) gesture classes and the second experiment has (10, 20, 30, 40) gesture classless.
Huang et al. (2023) Tumor subtype classification at WSI-level (4 Ep.) Incrementally learning tumor classes using 4 datasets (NSCLC, BRCA, RCC, ESCA) from TCGA [^1] https://www.cancer.gov/ccg/research/genome-sequencing/tcga project. Each dataset has 2 distinct tumor subtypes leading to a total 8 subtypes.
Xiao et al. (2023) Skin disease classification (18 Ep.) Incrementally learning 5 new skin disease classes in each episode
Li and Jha (2023) Disease classification on physiological signals (2 Ep.) 3 separate experiments were designed, each with 2 episodes. (a) using CovidDeep(Hassantabar et al. 2021), 1st episode has healthy and symptomatic patients and 2nd episode only asymptomatic patients, (b) using DiabDeep(Yin et al. 2019) dataset: 1st episode has healthy and Type-I diabetic patients and 2nd episode only Type-II diabetic patients, and (c) using MHDeep(Hassantabar et al. 2022) dataset: 1st episode has participants with healthy condition and major depressive disorder and 2nd episode has bipolar depressive disorder patients.
Table 5: List of various class incremental scenarios in literature (Continued)
Reference (year) Application (# Episodes) Description
Sun et al. (2023a) Multi-class classification on time-series signals (5 Ep., 10 Ep., 5 Ep., 4 Ep.) 4 separate experiments were designed as follows: (a) using Mit-BIH dataset (Goldberger et al. 2000) 4 episodes were created where 1st episode has 4 classes and the 4 new classes are added one by one in the next 4 episodes. (b) using FaceAll dataset (Dau et al. 2019), 10 episodes were created where 1st episode has 5 classes and the next 9 episodes introduce 9 new classes, one by one. (c) using Wave dataset (Dau et al. 2019) a total of 5 episodes were created where 1st episode has 12 lasses and the next 4 classes introduce new 12 classes, 3 classes each. (d) using Mit-BIH Long-Term ECG dataset (Goldberger et al. 2000) 4 episodes were created where 3 classes were in 1st episode and the next 3 episodes contained 3 new classes, 1 class in each.
Yang et al. (2023b) Skin disease classification on dermoscopic and clinical images: (4-20 Ep.) 2 datasets were used separately as follows: (a) Skin7 (Codella et al. 2018) containing a total of 7 classes is used in CIS where 1 and 2 classes are incrementally added leading to 6 and 4 episodes, respectively. (b) Skin40 dataset (Sun et al. 2016) containing a total of 40 classes is used in CIS where 2,5, and 10 classes are incrementally added leading to 20, 8, and 4 episodes, respectively.
Verma et al. (2023) Disease classification on (a) fundus (3 Ep.), (b) pathology images: (3 Ep.) (a) OCT dataset (Kermany et al. 2018) with total 4 classes used to create 3 episodes: episode-1 has two classes (Normal and Choroidal Neovascularization), episode-2 has 1 class (Diabetic Macular Edema) and episode-3 contains 1 class (Drusen). (b) PathMNIST dataset (Yang et al. 2023a) from the MedMNIST repository offers 9 classes for colon pathology. It is used to create 3 episodes with 3 classes each.
Ceccon et al. (2024a) Chest X-ray based disease classification (5 Ep.) 2 datasets (ChestX-ray14 (Wang et al. 2017) with 14 classes and CheXpert (Irvin et al. 2019) with 12 classes) used separately to create two experiments, each with 5 episodes.
Bayasi et al. (2024) (a) Skin lesion classification (2-3 Ep.), (b) blood cell classification (4 Ep.), (c) colon tissue classification (4 Ep.) (a) HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), and Derm7pt (Kawahara et al. 2018) datasets with total 7, 7, and 6 classes used to create 3, 3, and 2 episodes each with 2-3 non-overlapping skin disease classes, respectively, (b) PBS-HCB (Acevedo et al. 2020) dataset with a total of 8 classes used to create 4 episodes each with 2 non-overlapping blood cell classes, (c) NCT-CRC-HE (Kather et al. 2019) dataset with a total of 9 classes used to create 4 episodes each with 2-3 non-overlapping colon tissue classes.
Qazi et al. (2024) (a) Disease classification in histopathology images: (7 Ep.), (b) Skin lesion classification: (4 Ep.) (a) Using Path16 dataset 7 episodes were created as different sources of dataset with different classes from histopathology, (b) using skin8 dataset, a total of 4 episodes were created by adding 2 new classes per episode.
Zhu et al. (2024a) Histopathology WSI retrieval: (4 Ep.) Authors curate a sequence of 4 WSI datasets (NSCLC, RCC, BRCA and GAST) from TCGA. Each episode introduces a new dataset with 2 new classes of cancer, showing a CIS setting.
@@ 3.2 Class incremental scenarios
Class Incremental Scenario (CIS) refers to a situation where the aim is to adapt the model to accommodate novel classes from new data streams. Given that each batch of these unseen classes offers knowledge that largely differs from the previous batches, this scenario is anticipated to be the most challenging among continual learning scenarios. An instance of this challenge is incorporating new types of tissues into a tumor classification model or introducing novel cell types into a model designed for AI-assisted diagnosis in cell detection. In this scenario, there is only one incremental task that gets additions of new classes in subsequent episodes. Some works assume mutually exclusive classes, whereas some classes are overlapping.
@@
Zhang and Wang (2019b) Longitudinal MRI-based Alzheimer’s disease progression modeling (7 Ep.) Tasks are MR images belonging to different time points
Kaustaban et al. (2022) Histopathology data-based tumor classification (4 Ep.) 4 episodes were created in TIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes
Derakhshani et al. (2022) Disease classification (4 Ep.) 4 episodes were created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (MedMNIST repository (Yang et al. 2023a))
Bera et al. (2023) MRI-based binary segmentation (3 Ep.) 3 episodes were curated as binary segmentation application using MRI data on 3 organs (prostate, spleen, hippocampus): Promise12 (Litjens et al. 2014) (prostate) $\rightarrow$ MSD (NCI ) (spleen) $\rightarrow$ Drayd (Denovellis et al. 2021) (hippocampus)
Wu et al. (2024) Image super-resolution (4 Ep.) First 3 episodes are curated from IXI (IXI ) dataset referring to (PD, T1, and T2) weighted brain MRI, and the last episode refers to chest X-ray from Chest X-ray (Wang et al. 2017) dataset.
Li and Jha (2023) Disease classification on physiological signals (2-3 Ep.) A task incremental scenario with 2-3 different disease classification tasks were created using 2-3 datasets (CovidDeep(Hassantabar et al. 2021), DiabDeep(Yin et al. 2019), and MHDeep(Hassantabar et al. 2022))
Verma et al. (2024) Disease classification on (a) Fundus (2 Ep.), (b) pathology images (3 Ep.) (a) OCT (Kermany et al. 2018) with total 4 classes used to create 2 episodes, each having 2 classes (b) PathMNIST (Yang et al. 2023a) from the MedMNIST repository offers 9 classes for colon pathology. It is used to create 3 episodes, each with 3 classes.
Ye et al. (2024) Multi-modality representation learning (5 Ep.) 5 different medical data modalities including medical report, MRI, X-ray, CT, and histopathology data were learned in 5 episodes for SSL-based representation learning